A) An entrepreneur founded his company using $200,000 of his own money, issuing himself 200,000 shares of stock. An angel investor bought an additional 100,000 shares for $150,000. The entrepreneur now sells another 400,000 shares of stock to a venture capitalist for $2 million. What is the post-money valuation for the last round of funding in dollars?

B) In previous question, suppose the company intends to go public by selling 3,000,000 new shares. Moreover, assume the company has no debt but has an excess cash of $20 million, and the revenues for next year is estimated to be $25 million. Further assume thad the industry average forward EV-to-Sales multiple is 4, whereas the average trailing EV-to-Sales ratio is 5. What is the estimated IPO stock price?
Answer #1

A)

Initially, the entrepreneur invested $200,000 and issued 200,000 shares. This makes the value per share as $1

After an angel investor agrees to invest $150,000 and in turn is issued 100,000 shares. This makes the value per share as $1.5

Thus, the total value of the company = total number of shares * value per share = (200,000 + 100,000)* $1.5 = 300,000 * $1.5 = $450,000

Now, when the venture capitalist invests $2 million and is issued 400,000 shares in exchange. This values the company at $5/share

Thus, the total value of the company = Total number of shares * value per share

The post money valuation = (200,000 + 100,000 + 400,000) * $5 = 700,000 * $5 = $3.5 million

B)

The average forward EV-to-Sales ratio is 4 and the forward looking sales is $25 million

Thus EV = $25 million * 5 = $100 million

Now the total number of shares = 700,000 + 300,000 = 1 million

EV = Equity - Cash

Thus, Equity = EV + Cash = $100 million + $20 million = $120 million

Thus, the value per share = EV/Total number of shares = $120 million/1 million

Value per share = $120

Thus, the estimated IPO stock price should be $120